When 3.802 g of Na3 PO4 ·12H2O react with excess BaCl2 ·2H2O, how many moles of Ba3 (PO4 ) 2 would be produced?


1
Expert's answer
2021-01-04T03:49:28-0500

2Na3PO4·12H2O + 3BaCl2·2H2O → Ba3(PO4)2 + 6NaCl + 30H2O

M(Na3PO4·12H2O) = 380.124 g/mol

n(Na3PO4·12H2O) "= \frac{3.802}{380.124}" = 0.01 mol

n(Ba3(PO4)2) "= \frac{1}{2}" n(Na3PO4·12H2O) "= \frac{1}{2} \times 0.01" = 0.005 mol

Answer: 0.005 mol